# TockyPrep: Data Preprocessing Methods for Flow Cytometric Fluorescent Timer Analysis

### v1

@@ 1 Background
Fluorescent Timer proteins are distinguished by their unique ability to exhibit time-dependent changes in their emission spectra. A variety of these proteins have been developed, including DsRed-Timer [1], mCherry-derived monomeric variants such as Fast-FT, Medium-FT, and Slow-FT [2], and mRuby variants, known as mRubyFT [3]. These proteins have enabled detailed analyses of temporal dynamics in molecular and cellular events. Applications range from examining intracellular and membranous protein dynamics [2] to monitoring T cell and B cell activities in vivo [4], and assessing promoter activities [1] [5]. Furthermore, tandem fluorescent protein timers, pairing two fluorescent proteins with distinct expression and maturation kinetics, are employed in both cellular and in vivo studies. The combination of Superfolder GFP (sfGFP), a GFP variant known for its rapid fluorescence maturation, with a red fluorescent protein exhibiting more delayed fluorescence kinetics, such as mCherry or TagRFP, is a popular choice for these tandem Timer systems [6, 7]. The analysis of data from these studies typically involves sophisticated mathematical modeling to elucidate temporal changes in gene expression by exploiting the differing kinetics of the two fluorescent proteins [7].
Using the monomeric Fluorescent Timer protein Fast-FT, we have developed the Timer-of-cell-kinetics-and-activity (Tocky) system for single-cell flow cytometric analysis of cellular activities and transcription [4]. This approach has led to the creation of multiple Fluorescent Timer reporter mouse strains, such as Foxp3-Tocky and Nr4a3-Tocky, enabling detailed analysis of the temporal dynamics of Foxp3 transcription and transcriptional activities downstream of T cell receptor (TCR) signaling. These developments have established the proof-of-concept for the Tocky systems as valuable experimental methods[4][8]. However, the full implementation of Timer fluorescence data analysis has yet to be realized.
@@
Importantly, flow cytometric measurements of fluorescence rely on signal amplification [9], and the instrumental settings, especially the voltages used for amplification, can be significantly varied between users [10]. This variability complicates the consistent application of trigonometric transformations, underscoring the necessity for robust and practical data preprocessing methods. Consequently, the normalisation of immature and mature Timer fluorescence data becomes crucial in ensuring that the analysis of Timer blue and red is reliable and reproducible across different experimental setups.
In this paper, we introduce a new implementation of the Timer data analysis method in R. This implementation automates the preprocessing, normalisation, and trigonometric transformation steps essential for analysing Timer data from flow cytometry experiments, aiming to standardize the process and enhance reproducibility and accuracy.
@@ 2.1 Overview of Timer Data Analysis
    Timer fluorescence normalisation, which normalizes Timer blue and red fluorescence data.
@@
  4. 4.
    Data export for downstream analysis
@@ 2.3 Timer Fluorescence Normalization Using Negative Control Data
where $\max(B_{\text{log, gated\_neg}})$ is the maximum value and $\text{SD}(B_{\text{log, gated\_neg}})$ is the standard deviation of the log-transformed blue fluorescence of gated negative cells.
@@
where $\max(R_{\text{log, gated\_neg}})$ is the maximum value and $\text{SD}(R_{\text{log, gated\_neg}})$ is the standard deviation of the log-transformed red fluorescence of gated negative cells.
@@ 2.4 Timer Fluorescence Trigonometric Transformation
This angle effectively maps the maturation state of the fluorescent protein, with values ranging from $0^{\circ}$ (pure blue fluorescence) to $180^{\circ}$ (pure red fluorescence), reflecting the time elapsed since expression.
@@ 2.5 Nr4a3 Tocky Dataset
In this study, we assessed the effectiveness of Timer fluorescence normalization using flow cytometric data from antigen-stimulated T-cells, specifically obtained from Nr4a3-Tocky::OT-II double transgenic mice. These T-cells, expressing the OT-II transgenic T-cell receptor, were activated using their specific agonist peptide, Ova(323-339) at a concentration of 1 $\mu$M in the presence of antigen presenting cells. This setting provides continuous T-cell receptor signals to T-cells and induce Fluorescent Timer expression. Time course analysis was performed. This dataset was previously detailed in our earlier publication [4], where we demonstrated the initial application of the Tocky system in a biological context.
@@ 2.7 Software
The plots in Figures 7 and 9 were generated using the R packages ggplot2[11], gridExtra[12], and dplyr[13]. The package RColorBrewer[14] was used to generate heat colors. All other figures were generated by the TockyPrep package.
@@ 3.1 Fluorescent Timer Protein
Fluorescent Timer proteins have two types of fluorescence, and the different kinetics of these fluorescence levels are exploited to decipher temporal dynamics of biological systems. We use an mCherry mutant Fluorescent Timer, Fast-FT [2]. Similar to mCherry, Fast-FT emits blue fluorescence soon after translation. However, the chromophore spontaneously matures into the red form. Our previous experimental measurement of the maturation half-time was 4 hours [8]. The two different types of Timer fluorescence—Timer Blue fluorescence from the immature chromophore and Timer Red fluorescence from the mature chromophore—are excited by different lasers (402 nm laser and 561 nm laser, respectively) and analyzed and measured by different detectors (Figure 1A).
@@ 3.2 Overview of TockyPrep
Figure 1: Fluorescent Timer protein and workflow for Timer fluorescence data preprocessing using the TockyPrep package. (A) Diagram illustrating that Timer Blue fluorescence from the immature chromophore is excited by a 402 nm laser, and Timer Red fluorescence from the mature chromophore is excited by a 561 nm laser. These fluorescences are detected by different detectors during flow cytometry. (B) Overview of the TockyPrep workflow, showing data setup using prep_tocky and preprocessing steps executed by timer_transform, which include Timer Thresholding, Timer Fluorescence Normalization, and Trigonometric Transformation.
@@ 3.3 Workflow
Figure 2: The two central functions in the TockyPrep implementation, prep_tocky and timer_transform. The function prep_tocky sets up sample CSV files and identifies key variables, leading into timer_transform for executing comprehensive data preprocessing steps.
@@ 3.6 Timer Fluorescence Normalization
To address these challenges, the TockyPrep package introduces a robust function specifically designed for Timer fluorescence normalization. This function leverages autofluorescence levels to accurately estimate inherent signal biases. Moreover, it applies the standard deviation (SD) as a normalization factor to correct differences between Blue and Red fluorescence signals. By doing so, it aims to enhance the precision of Timer Angle calculations.
Figure 4: Timer Fluorescence Normalization. This figure illustrates the challenges in Timer fluorescence data analysis: (1) the absence of a calibration standard for Timer Blue, (2) acquisition settings that induce signal bias between Timer Blue and Red fluorescence, and (3) the resultant skew in Timer Angle estimation due to these biases. To address these issues, the TockyPrep package includes a function for Timer fluorescence normalization. This function utilizes autofluorescence levels to estimate inherent signal biases and employs standard deviation (SD) to normalize differences between Blue and Red fluorescence, thereby enhancing the accuracy of Timer Angle calculations.
@@ 4 Results
In this investigation, we analyzed the flow cytometric dataset that analyzed activated T cells obtained from Nr4a3-Tocky::OT-II double transgenic mice[4]. These T cells, which express a specific transgenic T-cell receptor (OT-II), were activated with 1 $\mu$M Ova(323-339) peptide to assess the dynamic transcriptional activity of Nr4a3 following antigen stimulation in a time course analysis [4]. This dataset offers opportunities to address how Tocky data preprocessing is applied.
@@ 4.2 Analyzing Activated T Cells from Nr4a3 Tocky Mice Using TockyPrep
The TockyPrep package includes a plot function, plot_tocky, dedicated to visualizing Timer fluorescence expression and demonstrating how Timer thresholds are applied to experimental samples (Figure 6). The effects of activation by Ova peptide are readily visible in conventional 2D plots for Timer fluorescence: Timer Blue fluorescence is quickly induced shortly after activation, followed by an increase in Timer Red fluorescence (Figure 6). Importantly, the quadrant gates displayed in the output from plot_tocky represent the Timer thresholds that have been set by timer_transform, underscoring the practical application of this visualization tool.
Notably, T-cells at the 0-hour time point include Timer Blue-Red+ cells, comprised of self-reactive T-cells that accumulate Timer Red fluorescence in vivo [16]. Since this accumulation occurs during antigen recognition in vivo and not from in vitro antigen stimulation, this time point is excluded from downstream analysis.
Figure 6: Visualization of Timer Thresholding Using plot_tocky. This output demonstrates the application of Timer fluorescence thresholds to experimental samples, highlighting the rapid induction of Timer Blue fluorescence following activation, followed by an increase in Timer Red fluorescence. Quadrant gates show the thresholds set by timer_transform.
Once T cells recognize an antigen, they receive T-cell receptor signaling and become activated, exhibiting increased cell size as measured by forward scatter (FSC) and elevated Timer expression. As expected, FSC significantly increases following activation (Figure 7A). Similarly, both Timer Blue and Red fluorescence intensify post-activation, as demonstrated by the analysis of mean fluorescence intensities (Figure 7B and C).
Figure 7: Changes in FSC and Timer Fluorescence After T-Cell Activation. (A) FSC measurements showing increased cell size after activation. (B) Timer Blue fluorescence increases after activation. (C) Timer Red fluorescence increases after activation. Note that the time point zero is excluded due to the accumulation of self-reactive T-cells that accumulate Timer red fluorescence.
After applying the timer_transform function, we utilized plot_tocky with the normalize = TRUE option to visualize the effects of Timer thresholding and normalization on Timer Blue and Red fluorescence (Figure 8). Importantly, the data displayed in this visualization reflect not only the normalization of Timer Blue and Red fluorescence but also the application of Timer thresholding, ensuring that all analyzed cells are Timer-positive.
Figure 8: Visualization of Timer Fluorescence Before and After Normalization. Comparison of Timer Blue and Red fluorescence with and without normalization using plot_tocky. The normalization process adjusts fluorescence intensities to correct for biases in data acquisition.
@@
Finally, we examined Timer Angle and Intensity using the function plot_angle, which produces 2D plots for Timer Angle and Intensity. Similar to plot_tocky, plot_angle can produce group-wise 2D plots. The Timer Angle-Intensity plots effectively show the progressive increase of Timer expression from low Timer Angle values, gradually accumulating cells with high Timer Angle and Intensity values.
Figure 10: Visualization of Timer Angle and Intensity After Normalization. 2D plots generated by plot_angle, showing the distribution of Timer Angle and Intensity in the experimental samples after applying Timer fluorescence normalization. The plots illustrate the progressive increase of Timer expression and accumulation of cells with higher Timer Angle and Intensity values.
@@ 5 Discussion
Our investigations identified the normalization of Timer fluorescence between Timer blue and red fluorescence was key for quantitative analysis of Timer maturation using the trigonometric transformation. Since the current method relies on autofluorescence data, ubiquitous adjustments through normalization using the standard deviations (SDs) of Timer Blue and Red fluorescence may not achieve the most optimized results for cells with extremely high expression. Still, our current investigations demonstrated the importance of Timer fluorescence normalization and the utility of the current method and implementation, providing reasonable and feasible solutions. Future studies can focus on the development of experimental standards to further enhance the normalization methods.
Our previous study utilized FSC data to correct size-dependent effects in Timer fluorescence [4], a method not implemented in the current development. Although traditional flow cytometry often shows higher background in the blue fluorescence channel, including Timer Blue fluorescence, the principle underlying FSC correction remains valid. However, in analyzing real datasets, we identified two significant issues. First, the FSC correction method depends on Timer-negative control samples, which typically lack activated cells with larger sizes. These cells are often removed during the quality control stage, complicating the determination of optimal FSC correction parameters across the full range of FSC and Timer fluorescence. Second, modern flow cytometers, especially spectral analyzers, show greater robustness against cell size effects. Consequently, we opted not to implement the FSC correction method in the current version of the package.
To date, various Fluorescent Timer protein reporter systems have been reported. The reporter systems using Fast-FT include Nr4a3-Tocky for the TCR signal downstream gene Nr4a3 [4], Foxp3-Tocky for the regulatory T cell-specific gene Foxp3 [4, 8], HIV-Tocky for the long terminal repeat (LTR) of HIV [17], Nur77-Tempo for another TCR signal downstream gene, Nur77 [18], and H2B, a histone gene [19]. In addition, other fluorescent Timer proteins like DsRed-E5 have been utilized to develop a transgenic reporter for the proglucagon gene Gcg [20]. Particularly, Nr4a3-Tocky has been used for the broadest range of applications including tumour immunology and immunotherapy [21][22] NKT cell biology[23], and autoimmune disease[4].

### v3

@@ 1 Background
Fluorescent Timer proteins are distinguished by their unique ability to exhibit time-dependent changes in their emission spectra. A variety of these proteins have been developed, including DsRed-Timer [1], mCherry-derived monomeric variants such as Fast-FT, Medium-FT, and Slow-FT [2], and mRuby variants, known as mRubyFT [3]. These proteins have enabled detailed analyses of temporal dynamics in molecular and cellular events. Applications range from examining intracellular and membranous protein dynamics [2] to monitoring T cell and B cell activities in vivo [4], and assessing promoter activities [1, 5]. Furthermore, tandem fluorescent protein timers, pairing two fluorescent proteins with distinct expression and maturation kinetics, are employed in both cellular and in vivo studies. The combination of Superfolder GFP (sfGFP), a GFP variant known for its rapid fluorescence maturation, with a red fluorescent protein exhibiting more delayed fluorescence kinetics, such as mCherry or TagRFP, is a popular choice for these tandem Timer systems [6, 7]. The analysis of data from these studies typically involves sophisticated mathematical modeling to elucidate temporal changes in gene expression by exploiting the differing kinetics of the two fluorescent proteins [7].
Using the monomeric Fluorescent Timer protein Fast-FT, we have developed the Timer-of-cell-kinetics-and-activity (Tocky) system for single-cell flow cytometric analysis of cellular activities and transcription [4]. This approach has resulted in the development of Fluorescent Timer reporter mouse strains: Foxp3-Tocky for analyzing the temporal dynamics of Foxp3 transcription and Nr4a3-Tocky for studying cellular responses over time following T cell receptor (TCR) signaling. These developments have established the proof-of-concept for the Tocky systems as valuable experimental methods[4][8]. However, the full implementation of Timer fluorescence data analysis has yet to be realized.
@@
Importantly, flow cytometric measurements of fluorescence rely on signal amplification [9], and the instrumental settings, especially the voltages used for amplification, can be significantly varied between users [10]. This variability complicates the consistent application of trigonometric transformations, underscoring the necessity for robust and practical data preprocessing methods. Consequently, the normalization of immature and mature Timer fluorescence data becomes crucial in ensuring that the analysis of Timer blue and red is reliable and reproducible across different experimental setups.
In this paper, we introduce a new R implementation of the Timer data analysis method. This implementation automates the data preprocessing methods, including Timer fluorescence normalization and trigonometric transformation, which are essential for analyzing Timer data from flow cytometry experiments. Additionally, tailored visualization techniques are included to ensure accurate data preprocessing. Collectively, these enhancements aim to standardize the process, enhancing both reproducibility and accuracy.
@@ 2.1 Overview of Timer Data Analysis
    Timer fluorescence normalization, which normalizes Timer blue and red fluorescence data.
@@
The preprocessed data, specifically Timer Angle and Timer Intensity, are subsequently used for downstream analysis.
@@ 2.3 Timer Fluorescence Normalization Using Negative Control Data
where $\max(B_{\text{log, gated\_neg}})$ is the maximum value and $\text{SD}(B_{\text{log, gated\_neg}})$ is the SD of the log-transformed blue fluorescence of gated negative cells.
@@
where $\max(R_{\text{log, gated\_neg}})$ is the maximum value and $\text{SD}(R_{\text{log, gated\_neg}})$ is the SD of the log-transformed red fluorescence of gated negative cells.
@@ 2.4 Timer Fluorescence Trigonometric Transformation
This angle effectively maps the maturation state of the fluorescent protein, with values ranging from $0^{\circ}$ (pure blue fluorescence) to $90^{\circ}$ (pure red fluorescence), reflecting the time elapsed since expression.
@@ 2.5 Nr4a3 Tocky Dataset
In this study, we assessed the effectiveness of Timer fluorescence normalization using flow cytometric data from antigen-stimulated T cells, specifically obtained from Nr4a3-Tocky::OT-II double transgenic mice. These T cells, expressing the OT-II transgenic TCR, were activated using their specific agonist peptide, Ova(323-339) at a concentration of 1 $\mu$M in the presence of antigen presenting cells. This setting provides continuous TCR signals to T cells and induce Fluorescent Timer expression. Time course analysis was performed. This dataset was previously detailed in our earlier publication [4], where we demonstrated the initial application of the Tocky system in a biological context.
@@ 2.7 Software
The plots in Figures 7 and 9 were generated using the R packages ggplot2[11], gridExtra[12], and dplyr[13]. The package RColorBrewer[14] was used to generate heat colors. All other figures were generated by functions implemented in the TockyPrep package. The interactive application for exploring Timer data preprocessing parameters was developed using the Shiny framework in R [15].
@@ 3.1 Fluorescent Timer Protein
Fluorescent Timer proteins have two types of fluorescence, and the different kinetics of these fluorescence levels are exploited to decipher temporal dynamics of biological systems. We use an mCherry mutant Fluorescent Timer, Fast-FT [2]. Similar to mCherry, Fast-FT emits blue fluorescence soon after translation. However, the chromophore spontaneously matures into the red form. Our previous experimental measurement of the maturation half-time was 4 hours [8]. The two different types of Timer fluorescence—Timer Blue fluorescence from the immature chromophore and Timer Red fluorescence from the mature chromophore—are excited by different lasers (405 nm laser and 561 nm laser, respectively) and analyzed and measured by different detectors (Figure 1A).
@@ 3.2 Overview of TockyPrep
Figure 1: Fluorescent Timer protein and workflow for Timer fluorescence data preprocessing using the TockyPrep package. (A) Diagram illustrating that Timer Blue fluorescence from the immature chromophore is excited by a 405 nm laser, and Timer Red fluorescence from the mature chromophore is excited by a 561 nm laser. These fluorescences are detected by different detectors during flow cytometric analysis. (B) Overview of the TockyPrep workflow, showing data setup using prep_tocky and preprocessing steps executed by timer_transform, which include Timer Thresholding, Timer Fluorescence Normalization, and Trigonometric Transformation.
@@ 3.3 Workflow
Figure 2: The two central functions in the TockyPrep implementation, prep_tocky and timer_transform. The function prep_tocky determines the sample CSV files for processing and identifies key variables, enabling timer_transform to execute data preprocessing steps in a batch manner.
@@ 3.6 Timer Fluorescence Normalization
To address these challenges, the TockyPrep package introduces a robust function specifically designed for Timer fluorescence normalization. This function leverages autofluorescence levels to accurately estimate inherent signal biases. Moreover, it applies MAD by default as the normalization factor. These adjustments correct disparities between Timer Blue and Red fluorescence signals, thereby enhancing the precision of Timer Angle calculations.
Figure 4: Timer Fluorescence Normalization. This figure illustrates the challenges in Timer fluorescence data analysis: (1) the absence of a calibration standard for Timer Blue, (2) acquisition settings that induce signal bias between Timer Blue and Red fluorescence, and (3) the resultant skew in Timer Angle estimation due to these biases. To address these issues, the TockyPrep package includes a function for Timer fluorescence normalization. This function utilizes autofluorescence levels to estimate inherent signal biases and employs MAD to normalize differences between Blue and Red fluorescence, thereby enhancing the accuracy of Timer Angle calculations.
@@ 4 Results
In this investigation, we analyzed the flow cytometric dataset that analyzed activated T cells obtained from Nr4a3-Tocky::OT-II double transgenic mice[4]. These T cells, which express the transgenic TCR OT-II, were activated with 1 $\mu$M Ova(323-339) peptide to assess the dynamic transcriptional activity of Nr4a3 following antigen stimulation in a time course analysis [4]. This dataset offers opportunities to address how Tocky data preprocessing is applied.
@@ 4.2 Analyzing Activated T Cells from Nr4a3 Tocky Mice Using TockyPrep
The TockyPrep package includes a key plot function, plot_tocky, which visualizes Timer fluorescence variables to understand the effects of thresholding, normalization, and transformation. Using the option plot_mode = "Timer fluorescence", plot_tocky shows raw Timer fluorescence, demonstrating how Timer thresholds are applied to experimental samples (Figure 6). In Figure 6, dynamic changes of Timer profiles upon TCR stimulation are thus visualized in conventional 2D plots for Timer fluorescence: Timer Blue fluorescence is quickly induced shortly after activation, followed by an increase in Timer Red fluorescence. Importantly, the quadrant gates displayed in the output from plot_tocky represent the Timer thresholds that have been set by timer_transform, assisting the quality check of Timer data preprocessing.
Notably, T cells at the 0-hour time point include Timer Blue-Red+ cells, comprised of self-reactive T cells that accumulate Timer Red fluorescence in vivo [17]. Since this accumulation occurs during antigen recognition in vivo and not from in vitro antigen stimulation, this time point is excluded from downstream analysis.
Figure 6: Visualization of Timer Thresholding Using plot_tocky. This figure shows Timer Blue and Red fluorescence of T cells before normalization. T cells from Nr4a3-Tocky::OT-II double transgenic mice were activated by Ova peptide and analyzed at the indicated time points [4]. Quadrant gates illustrate the thresholds for Timer Blue and Red fluorescence set by timer_transform.
Once T cells recognize an antigen, they receive TCR signaling and become activated, exhibiting increased cell size as measured by forward scatter (FSC) and elevated Timer expression. As expected, FSC significantly increases following activation (Figure 7A). Similarly, both Timer Blue and Red fluorescence intensify post-activation, as demonstrated by the analysis of mean fluorescence intensities (Figure 7B and C).
Figure 7: Changes in FSC and Timer Fluorescence After T-Cell Activation. (A) FSC measurements showing increased cell size after activation. (B) Timer Blue fluorescence increases after activation. (C) Timer Red fluorescence increases after activation. Note that the time point zero is excluded due to the accumulation of self-reactive T cells that accumulate Timer red fluorescence.
After applying the timer_transform function, we utilized plot_tocky with the option plot_mode = "Normalized Timer fluorescence" to visualize the effects of Timer thresholding and normalization on Timer Blue and Red fluorescence (Figure 8). Importantly, the data displayed in this visualization reflect not only the normalization of Timer Blue and Red fluorescence but also the application of Timer thresholding, ensuring that all analyzed cells are Timer-positive.
Figure 8: Visualization of Normalized Timer Fluorescence Using plot_tocky. This figure displays the normalized Timer Blue and Red fluorescence, visualized using plot_tocky. The data shown here are the same as those in Figure 7, but presented with the option plot_mode = "Normalized Timer fluorescence".
@@
Finally, we examined Timer Angle and Intensity using the function plot_tocky with the option plot_mode = "Timer Angle and Intensity.", which produces group-wise 2D plots for Timer Angle and Intensity. The Timer Angle-Intensity plots effectively show the progressive increase of Timer expression from low Timer Angle values, gradually accumulating cells with high Timer Angle and Intensity values (Figure 10).
Figure 10: Visualization of Timer Angle and Intensity Using plot_tocky. Two-dimensional plots showing Timer Angle and Intensity were generated using the plot_tocky function with the option plot_mode = "Timer Angle and Intensity.". These plots display data that has been normalized and transformed into polar coordinates.
4.3 Exploring Timer Data Transformation Parameters
By using the function explore_timer_transform, the parameter space can be explored by adjusting parameters for Timer fluorescence normalization and transformation. Executing explore_timer_transform launches an interactive HTML application, enabling users to experiment with changing Timer fluorescence thresholds and selecting different normalization methods on their flow cytometric Timer data (Figure 11).
Figure 11: Exploration of Timer Data Transformation Parameters Using explore_timer_transform. This figure displays a screenshot of the interactive application outputs. The sidebars enable users to set thresholds for Timer Blue and Timer Red fluorescence and to choose a normalization method. The plot panels include 2D plots for raw Timer fluorescence (Upper Left), normalized Timer fluorescence (Upper Right), and Timer Angle and Timer Intensity (Lower Left) using all concatenated expression data. Additionally, a 2D plot of negative control data for raw Timer fluorescence is shown, with Timer thresholds determined by the sidebar settings.
@@ 5 Discussion
Our investigations identified the normalization of Timer fluorescence between Timer blue and red fluorescence was key for quantitative analysis of Timer maturation using the trigonometric transformation. Since the current method relies on autofluorescence data, ubiquitous adjustments through normalization using the MADs or SDs of Timer Blue and Red fluorescence may not achieve the most optimized results for cells with extremely high expression. Still, our current investigations demonstrated the importance of Timer fluorescence normalization and the utility of the current method and implementation, providing reasonable and feasible solutions. Future studies can focus on the development of experimental standards to further enhance the normalization methods.
Our previous study utilized FSC data to correct size-dependent effects in Timer fluorescence [4], a method not implemented in the current development. Since traditional flow cytometry often shows higher background in the blue fluorescence channel, including Timer Blue fluorescence, the principle underlying FSC correction remains valid. However, in analyzing real datasets, we identified two significant issues. First, the FSC correction method depends on Timer-negative control samples, which typically lack activated cells with larger sizes. These cells are often removed during the quality control stage, complicating the determination of optimal FSC correction parameters across the full range of FSC and Timer fluorescence. Second, modern flow cytometers, especially spectral analyzers, show greater robustness against cell size effects. Consequently, we opted not to implement the FSC correction method in the current version of the package.
To date, various Fluorescent Timer protein reporter systems have been reported. The initial murine reporter systems utilizing Fast-FT were Nr4a3-Tocky, designed to analyze the TCR signal downstream gene Nr4a3, and Foxp3-Tocky, developed for the regulatory T cell-specific gene Foxp3 [4, 8]. These systems pioneered the use of Fast-FT in analyzing transcriptional and cellular dynamics across cellular biology and immunological studies [16]. Subsequent developments have expanded to include HIV-Tocky for the long terminal repeat (LTR) of HIV [18], Nur77-Tempo for analyzing TCR signal downstream Nur77 (Nr4a1), another member of the Nr4a family, similar to Nr4a3-Tocky [19], and H2B-FT for a histone gene [20]. In addition, other fluorescent Timer proteins like DsRed-E5 have been utilized to develop a transgenic reporter for the proglucagon gene Gcg [21]. Notably, Nr4a3-Tocky has been the most frequently used for the broadest range of applications including tumour immunology and immunotherapy [22][23] NKT cell biology[24], and autoimmune disease[4].